OpenStudy (anonymous):

What is the solution to the rational equation x over x squared minus 9 plus 1 over x minus 3 equals 1 over 4x minus 12

OpenStudy (anonymous):

\[\frac{ x }{ x ^{2}-9 }+\frac{ 1 }{ x+3 } = \frac{ 1 }{ 4x-12 }\] Is this what it looks like?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

...

OpenStudy (anonymous):

Well you need to combine the left hand side of the equation into one denominator.

OpenStudy (anonymous):

Then you need to get rid of both the left hand side denominator and right hand side denominator by multiplying by each one on both sides. Then it should be easier to solve using the quadratic formula.